Grade 12th passMechanics

Work done in increasing the length of massless spring by 1cm from natural length is 10J.work done in increasing the length further by 1cm is
1- 10J
2- 30J
3-50J
4-80J

Ashish kumar
8 Years agoGrade 12th pass

1 Answer

Saketh Vaddamani
8 Years ago
30J is the answer because,
Let the spring constant be K then,
1/2×K×1=10
To increase length by further 1cm the work done is
1/2×K×{22-12}=1/2×K×3=30J